| 
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
Sample Environment
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
 | 
| 
SASE1 Status and SASE3 Readiness
 | 
|   Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  | 
Not logged in | 
  | 
   | 
| 
 | 
| 
Message ID: 564
    Entry time: 17 Sep 2018, 14:30
    In reply to: 563
 |  
| Author: | 
Dennis Goeries  | 
 | Location: | 
General  | 
 | Severity: | 
Information  | 
 | Subject: | 
Re: Status report after dispatch meeting 09.30  | 
 
  |  
 
 | 
The GUI crashes reported by FXE will be fixed with the hotfix release coming ~tomorrow. 
It is related to very high or low values in plotting widgets scratching the maximum floating point number (1e308 ~ infinity). 
-- Wajid & Dennis 
  
	
		
			| Adriano Violante wrote: | 
		 
		
			| 
			 SPB 
			
				- 
				
Bug in the AGIPD cooling macro (setpoint resets itself to -59 C)  -  Jola knows about the problem. AS: Macro will be fixed. 
				 
				- 
				
AGIPD dark calibration: when new darks were injected with calling Steffen's script "xfel-calibrate",  error message "Disk quota exceeded" popped up caused by a GPFS-related issue -> issue was solved. In general the script was not tested (scratch space was used!). One should not release script without ITDM knowing and haveing thested it before.  
				 
			 
			FXE 
			
				- Commissioning going on in the last days. Not able to get 14 keV beam for next users experiments. They will stick with 9.3 keV.
 
				- Karabo stability issue (GUI) reported in the commissionig logbook.
 
			 
			SQS 
			
				- Problem with the digitizers bandwidth has now been solved.
 
				- Loop 6 running w/o smaract stages.
 
				- Flow switches issues solved . Rewiring of PLC needed
 
				- Pfeiifer flow valves ready to be stested this week.
 
				- Loop assignment changed. Change request handed in. As soon as it is approved it will be clear what loop to bring up next.
 
			 
			​SCS 
			
				- Loop 7 is running. Needs to be connected to Karabo.
 
				- Loop 8 running too.
 
				- Readiness of loop 1 and 4 ==> Interlock conditions need to be implemented. 
 
				- It is unclear whether the homing capability has been implemented already and can be used. Discussion ongoing. The use case has to be clarified.
 
				- Network storage for the control PCs in the control room is requested. Scope definition and discussions with CAS and ITDM will be done offline.
 
			 
			MID 
			
				- Loop 1 deployed and technical commissioning of the OPT hutch is ongoing. This week Slit and ATT will be commissioned. No big issues so far.
 
				- Loop 4 being prepared by AE. Will be ready next week.
 
			 
			HED 
			
				- Work in tunnel ongoing
 
				- HRMono- ready to start grouting. Mono will be here next week.
 
			 
			DET 
			
				- FXE/SPB GPUs crash issue  - to be checked with ITDM - ITDM was not contacted.
 
			 
			ITDM 
			
				- DAQ was upgraded to the user friendly version in SASE3. After having discussed with FXE in the meeting, it was decided to install it also in FXE. 
 
				- DAQ training requested by SQS
 
			 
			CAS 
			
				- Karabo 2.2.5. patch being prepared. Will solve minor bugs and the crash due to empty data files.
 
				- 28 Mpot update requests were received by CY. Need to be planned and coordinated with H.Santos.
 
			 
			  
			 | 
		 
	
 
   |